# Pruning Spurious Subgraphs for Graph Out-of-Distribtuion Generalization

### v3

@@ 4 Proposed Method
[…]
@@
Here,$\epsilon$ is a value close to zero, $p_{ij}$ denotes the normalized probability of the edge $e_{ij}$, and $\mathcal{E}_{s}$ is the lowest $K\%$ of edges among all estimated edge weights $w_{ij}\in\mathcal{E}$ by the subgraph selector $t(\cdot)$.
@@ 5 Theoretical Analysis
$L(\theta;D)$ $:=\frac{1}{n}\sum_{(x_{i},x_{j},y,G)\sim D}l((x_{i},x_{j},y,G);\theta),\forall e_{ij}\in\mathcal{E}.$ (7)
$L(\theta;S)$ $:=\frac{1}{m}\sum_{(x_{i},x_{j},y,G)\sim S}l((x_{i},x_{j},y,G);\theta),\forall e_{ij}\in\mathcal{E}.$
@@ 6 Related Work
ERM 68.66$\pm$4.25 51.74$\pm$2.88 69.58$\pm$2.51 59.94$\pm$2.37 62.77$\pm$2.14 61.03$\pm$1.88 64.93$\pm$6.25 68.10$\pm$1.68 78.29$\pm$3.76
IRM 70.65$\pm$4.17 51.41$\pm$3.78 67.97$\pm$1.84 59.00$\pm$2.92 63.96$\pm$3.21 62.47$\pm$1.15 72.27$\pm$3.41 67.22$\pm$1.15 77.56$\pm$2.48
GroupDRO 68.24$\pm$8.92 51.95$\pm$5.86 70.64$\pm$2.57 58.98$\pm$2.16 64.13$\pm$1.81 59.06$\pm$1.50 70.52$\pm$3.38 66.47$\pm$2.39 79.27$\pm$2.43
VREx 71.47$\pm$6.69 52.67$\pm$5.54 70.77$\pm$2.84 58.53$\pm$2.88 64.23$\pm$1.76 63.54$\pm$1.03 68.23$\pm$3.19 68.74$\pm$1.03 78.76$\pm$2.37
DropEdge 45.08$\pm$4.46 45.63$\pm$4.61 70.78$\pm$1.38 58.53$\pm$1.26 63.91$\pm$2.56 61.93$\pm$1.41 73.79$\pm$4.06 66.49$\pm$1.55 78.32$\pm$3.44
$G$-Mixup 59.66$\pm$7.03 52.81$\pm$6.73 70.01$\pm$2.52 59.34$\pm$2.43 61.90$\pm$2.08 61.06$\pm$1.74 69.28$\pm$1.36 67.44$\pm$1.62 78.55$\pm$4.16
FLAG 61.12$\pm$5.39 51.66$\pm$4.14 68.45$\pm$2.30 60.59$\pm$2.95 64.98$\pm$0.87 64.28$\pm$0.54 74.91$\pm$1.18 67.69$\pm$2.36 79.26$\pm$2.26
LiSA 54.59$\pm$4.81 53.46$\pm$3.41 70.38$\pm$1.45 52.36$\pm$3.73 62.60$\pm$3.62 60.96$\pm$1.07 69.73$\pm$0.62 68.11$\pm$0.52 78.62$\pm$3.74
DIR 62.07$\pm$8.75 52.27$\pm$4.56 68.07$\pm$2.29 58.08$\pm$2.31 63.91$\pm$2.92 61.91$\pm$3.92 66.13$\pm$3.01 66.86$\pm$2.25 76.40$\pm$4.43
DisC 51.08$\pm$3.08 50.39$\pm$1.15 68.07$\pm$1.75 58.76$\pm$0.91 59.10$\pm$5.69 57.64$\pm$1.57 61.94$\pm$7.76 67.12$\pm$2.11 56.59$\pm$10.09
CAL 65.63$\pm$4.29 51.18$\pm$5.60 67.37$\pm$3.61 57.95$\pm$2.24 65.03$\pm$1.12 60.92$\pm$2.02 74.93$\pm$5.12 68.06$\pm$2.60 79.50$\pm$4.81
GREA 56.74$\pm$9.23 54.13$\pm$10.02 67.79$\pm$2.56 60.71$\pm$2.20 64.67$\pm$1.43 62.17$\pm$1.78 71.12$\pm$1.87 69.72$\pm$1.66 77.34$\pm$3.52
GSAT 60.42$\pm$9.32 53.20$\pm$8.35 68.66$\pm$1.35 58.06$\pm$1.98 65.12$\pm$1.07 61.90$\pm$2.12 74.77$\pm$4.31 66.78$\pm$1.45 75.63$\pm$3.83
CIGA 68.71$\pm$10.9 49.14$\pm$8.34 69.40$\pm$2.39 59.55$\pm$2.56 65.42$\pm$1.53 64.47$\pm$0.73 74.94$\pm$1.91 64.92$\pm$2.09 65.98$\pm$3.31
AIA 72.91$\pm$5.62 55.85$\pm$7.98 71.15$\pm$1.81 61.64$\pm$3.37 64.71$\pm$0.50 63.43$\pm$1.35 76.01$\pm$1.18 70.79$\pm$1.53 81.03$\pm$5.15
PrunE 91.48*$\pm$0.40 66.53*$\pm$8.55 71.84*$\pm$0.61 64.99*$\pm$1.63 67.56*$\pm$0.34 65.46*$\pm$0.88 78.01*$\pm$0.42 70.32$\pm$1.73 81.59$\pm$5.35
@@ 7.4 Hyper-parameter Sensitivity
$\epsilon=0.01$ 91.63$\pm$0.73 60.38$\pm$8.35 77.76$\pm$1.11
$\epsilon=0.1$ 88.14$\pm$0.67 62.38$\pm$10.76 76.65$\pm$1.92
$\epsilon=0.3$ 80.93$\pm$4.33 50.65$\pm$4.95 76.07$\pm$2.65
$\epsilon=0.5$ 74.52$\pm$19.89 50.28$\pm$8.35 75.93±$\pm 1.27$
$\epsilon=\frac{1}{|\mathcal{E}|}$ 91.48$\pm$0.40 66.53$\pm$8.55 78.01$\pm$0.42
@@ F.1 Proof of Proposition 1
$\mathcal{L}_{GT}=-\mathbb{E}_{\mathcal{G}}\left[\log\mathbb{P}(Y\mid f(\widetilde{G}))\right],$ (12)
@@
$L(\theta;D)$ $:=\frac{1}{n}\sum_{\left(x_{i},x_{j},y,G\right)\sim D}l\left(\left(x_{i},x_{j},y,G\right);\theta\right),$ (14)
$L(\theta;S)$ $:=\frac{1}{m}\sum_{\left(x_{i},x_{j},y,G\right)\sim S}l\left(\left(x_{i},x_{j},y,G\right);\theta\right),$
@@ Appendix G More Discussion on Pruning-based Learning Paradigm
{tblr}
width = colspec = Q[110]Q[92]Q[106]Q[144]Q[106]Q[108]Q[106]Q[144], rowodd = c, row4 = c, cell12 = c=20.198, cell14 = c=20.25, cell16 = c=30.358, cell22 = c, cell23 = c, cell24 = c, cell25 = c, cell26 = c, cell27 = c, cell28 = c, hline1,5 = -0.08em, hline2 = 2-8lr, hline3 = -, & Motif GOODHIV EC50
@@ Appendix I The Pitfall of Directly Identifying Edges in $G_{c}$
SPMotif [52]. In SPMotif datasets [52], each graph comprises a combination of invariant and spurious subgraphs. The spurious subgraphs include three structures (Tree, Ladder, and Wheel), while the invariant subgraphs consist of Cycle, House, and Crane. The task for a model is to determine which one of the three motifs (Cycle, House, and Crane) is present in a graph. A controllable distribution shift can be achieved via a pre-defined parameter $b$. This parameter manipulates the spurious correlation between the spurious subgraph $G_{s}$ and the ground-truth label $Y$, which depends solely on the invariant subgraph $G_{c}$. Specifically, given the predefined bias $b$, the probability of a specific motif (e.g., House) and a specific base graph (Tree) will co-occur is $b$ while for the others is $(1-b)/2$ (e.g., House-Ladder, House-Wheel). When $b=\frac{1}{3}$, the invariant subgraph is equally correlated to the three spurious subgraphs in the dataset.
@@
Training and Validation. By default, we use Adam optimizer [21] with a learning rate of $31e-$ 3 and a batch size of 64 for all experiments. For DrugOOD, GOOD-Motif and GOOD-HIV datasets, our method is pretrained for 10 epochs with ERM, and for other datasets, we do not use ERM pretraining. We employ an early stopping of 10 epochs according to the validation performance for DrugOOD datasets and GOOD-Motif datasets, and do not employ early stopping for other datasets. Test accuracy or ROC-AUC is obtained according to the best validation performance for all experiments. All experiments are run with 4 different random seeds, the mean and standard deviation are reported using the 4 runs of experiments.
@@
For IRM and VREx, the weight of the penalty loss is searched over $\{1e-2,1e-1,1,1e1\}$. For GroupDRO, the step size is searched over $\{1.0,1e-1,1e-2\}$. The causal subgraph ratio for DIR is searched across $\{1e-2,1e-1,0.2,0.4,0.6\}$. For DropEdge, the edge masking ratio is seached over: $\{0.1,0.2,0.3\}$. For GREA, the weight of the penalty loss is tuned over $\{1e-2,1e-1,1.0\}$, and the causal subgraph size ratio is tuned over $\{0.05,0.1,0.2,0.3,0.5\}$. For GSAT, the causal graph size ratio is searched over $\{0.3,0.5,0.7\}$. For CIGA, the contrastive loss and hinge loss weights are searched over $\{0.5,1.0,2.0,4.0,8.0\}$. For DisC, we search over $q$ in the GCE loss: $\{0.5,0.7,0.9\}$. For LiSA, the loss penalty weights are searched over:$\{1,1e-1,1e-2,1e-3\}$. For $\mathcal{G}$-Mixup, the augmented ratio is tuned over $\{0.15,0.25,0.5\}$. For FLAG, the ascending steps are set to 3 as recommended in the paper, and the step size is searched over $\{1e-3,1e-2,1e-1\}$. For AIA, the stable feature ratio is searched over $\{0.1,0.3,0.5,0.7,0.9\}$, and the adversarial penalty weight is searched over $\{0.01,0.1,0.2,0.5,1.0,3.0,5.0\}$.
Hyperparameter search for PrunE. For PrunE, the edge budget $\eta$ in $\mathcal{L}_{e}$ is searched over: $\small\{0.5,0.75,0.85\small\}$; $K$ for the $K\%$ edges with lowest probability score in $\mathcal{L}_{s}$ is searched over:$\small\{50,70,90\small\}$; $\lambda_{1}$, $\lambda_{2}$ for balancing $\mathcal{L}_{e}$ and $\mathcal{L}_{s}$ are searched over: $\small\{10,40\small\}$ and $\small\{1e-1,1e-2,1e-3\small\}$ respectively. The encoder of subgraph selector $t(\cdot)$ is searched over $\{GIN,GCN\}$, with the number of layers: $\{2,3\}$.
@@
KMotif-base Motif-size

### v4

# Pruning Spurious Subgraphs for Graph Out-of-Distribution Generalization

@@ 4 Proposed Method
here $\|$ denotes the concatenation operator. To ensure the sampling process from $w_{ij}$ is differentiable and facilitate gradient-based optimization, we leverage the Gumbel-Softmax reparameterization trick bengio2013estimating ; maddison2016concrete , which is applied as follows:
@@
[…]
@@ 5 Theoretical Analysis
$L(\theta;D)$ $=\frac{1}{n}\sum_{(x_{i},x_{j},y,G)\sim D}l((x_{i},x_{j},y,G);\theta),\forall e_{ij}\in\mathcal{E}.$ (7)
$L(\theta;S)$ $=\frac{1}{m}\sum_{(x_{i},x_{j},y,G)\sim S}l((x_{i},x_{j},y,G);\theta),\forall e_{ij}\in\mathcal{E}.$
@@ 6 Related Work
[…]
@@ 7.4 Hyper-parameter Sensitivity
[…]
@@ F.1 Proof of Proposition 1
[…]
@@
$L(\theta;D)$ $=\frac{1}{n}\sum_{\left(x_{i},x_{j},y,G\right)\sim D}l\left(\left(x_{i},x_{j},y,G\right);\theta\right),$ (14)
$L(\theta;S)$ $=\frac{1}{m}\sum_{\left(x_{i},x_{j},y,G\right)\sim S}l\left(\left(x_{i},x_{j},y,G\right);\theta\right),$
@@ Appendix G More Discussion on Pruning-based Learning Paradigm
Motif GOODHIV EC50
@@ Appendix I The Pitfall of Directly Identifying Edges in $G_{c}$
SPMotif [52]. In SPMotif datasets [52], each graph comprises a combination of invariant and spurious subgraphs. The spurious subgraphs include three structures (Tree, Ladder, and Wheel), while the invariant subgraphs consist of Cycle, House, and Crane. The task for a model is to determine which one of the three motifs (Cycle, House, and Crane) is present in a graph. A controllable distribution shift can be achieved via a pre-defined parameter $b$. This parameter manipulates the spurious correlation between the spurious subgraph $G_{s}$ and the ground-truth label $Y$, which depends solely on the invariant subgraph $G_{c}$. Specifically, given the predefined bias $b$, the probability of a specific motif (e.g., House) and a specific base graph (Tree) will co-occur is $b$ while for the others is (1b)/2(1-b)/2 (e.g., House-Ladder, House-Wheel). When $b=\frac{1}{3}$, the invariant subgraph is equally correlated to the three spurious subgraphs in the dataset.
@@
Training and Validation. By default, we use Adam optimizer [21] with a learning rate of 1e31e-3 and a batch size of 64 for all experiments. For DrugOOD, GOOD-Motif and GOOD-HIV datasets, our method is pretrained for 10 epochs with ERM, and for other datasets, we do not use ERM pretraining. We employ an early stopping of 10 epochs according to the validation performance for DrugOOD datasets and GOOD-Motif datasets, and do not employ early stopping for other datasets. Test accuracy or ROC-AUC is obtained according to the best validation performance for all experiments. All experiments are run with 4 different random seeds, the mean and standard deviation are reported using the 4 runs of experiments.
@@
[…]
@@
K% Motif-base Motif-size